OpenStudy (anonymous):

Please check my work before i submit. I'm falling behind with 3 days left. I will fan , award a medal and write a detailed testimonial.

OpenStudy (anonymous):

details are always nice what is the question?

OpenStudy (anonymous):

this is one of 28 questions i have :(

OpenStudy (anonymous):

that would be wrong because you need a minus sign

OpenStudy (anonymous):

divide by \(180\) and multiply by \(\pi\) i.e. stick a \(\pi\) next to it

OpenStudy (anonymous):

so i have the sign wrong ?

OpenStudy (anonymous):

\[-\frac{54}{180}=-\frac{3}{10}\]

OpenStudy (jhannybean):

Simply just \[-54 \cdot \frac{\pi}{180}\]

OpenStudy (anonymous):

i.e reduce, then put a \(\pi\) there get \[-\frac{3\pi}{10}\]

OpenStudy (jhannybean):

I know `/circ` makes little circles but OS needs to have a fluttering degree symbol.

OpenStudy (jhannybean):

@satellite73

OpenStudy (anonymous):

yeah i haven't figured that out yet \[30^o\] looks ugly

OpenStudy (jhannybean):

\[-54^\circ\]

OpenStudy (anonymous):

\[30^{\circ}\] is not worth the bother

OpenStudy (anonymous):

so i see i miss calculated a great deal :(

OpenStudy (anonymous):

yes, you did cheat and use a calculator to reduce, it will make your life a lot easier

OpenStudy (anonymous):

okay thanks , this is my next answer

OpenStudy (jhannybean):

You mean your next question?

OpenStudy (anonymous):

Im having my answers checked before i submit my work into the system @Jhannybean

OpenStudy (anonymous):

think that is a mistake too

OpenStudy (anonymous):

the amplitude is 2 for sure, and the period is \(\pi\) but the phase shift i believe no not right

OpenStudy (anonymous):

it was explained to me this way i thought it was right :( @satellite73

OpenStudy (jhannybean):

Fortrig functions: \( F(t) = Af(Bt – C) + D\)

OpenStudy (anonymous):

here is the deal there is some formula to use, but i never remember it set \[2x-\pi=0\] solve for \(x\) you get \[2x=\pi\\ x=\frac{\pi}{2}\] so that last one has a phase shift of \(\frac{\pi}{2}\)

OpenStudy (anonymous):

so it's just pi when it's simplified

OpenStudy (jhannybean):

What do you mean by that?

OpenStudy (anonymous):

simplifying 2pi /pi would make the phase shift just pi . i believe

OpenStudy (anonymous):

the phase pellet of the one you chose is \(\frac{\pi}{4}\) the phase shift of D is \(\frac{\pi}{2}\)

OpenStudy (jhannybean):

phase pellet?lmao

OpenStudy (anonymous):

ohhhh so my answer was backwards ... idk im starting to feel confused

OpenStudy (jhannybean):

Well first off, why would you think the Amplitude is -2 when they told you it was 2? That would help you eliminate some of your answer choices.

OpenStudy (anonymous):

okay i see now . my mistake

OpenStudy (anonymous):